# Detecting Out-of-distribution through the Lens of Neural Collapse

### v5

@@ 5 Related Work
OOD Detection Extensive research has focused on developing OOD detection algorithms. One line of work is post-hoc and builds upon pre-trained models. For example, Hendrycks et al. (2019); Liang et al. (2018); Liu et al. (2020); Sun et al. (2021), and Sun and Li (2022) design OOD score over the output space of a classifier. Meanwhile, Lee et al. (2018) and Sun et al. (2022) measure OOD-ness from the perspective of ID clustering in feature space. Our work extends the observation that ID features tend to cluster from the perspective of Neural Collapse. While existing work is more focused are certain classification tasks than others, our proposed OOD detector is tested to be highly versatile.
Another line of work explores the regularization of OOD detection in training. For example, DeVries and Taylor (2018) and Hsu et al. (2020) propose OOD-specific architecture whereas Huang and Li (2021) and Wei et al. (2022) design OOD-specific training loss. In particular, Tack et al. (2020) brings attention to representation learning for OOD detection and proposes an OOD-specific contrastive learning scheme. Our work does not belong to this school of thought and is not restricted to specific training schemes or architecture.
Neural Collapse Neural Collapse was first observed in Papyan et al. (2020). During Neural Collapse, the penultimate layer features collapse to class means, the class means and the classifier collapses to a simplex equiangular tight framework, and the classifier simplifies to adopt the nearest class-mean decision rule. Further work provides theoretical justification for the emergence of Neural Collapse (Han et al., 2021; Mixon et al., 2020; Zhou et al., 2022; Zhu et al., 2021). In addition, Zhu et al. (2021) derives an efficient training algorithm drawing inspiration from Neural Collapse. Our concurrent work, Ammar et al. (2023), also leverages insights from Neural Collapse for OOD detection. However, they tackle from the subspace perspective and largely overlook class-specific information revealed by Neural Collapse, which is essential for our work.
@@ 6 Conclusion
This work leverages insights from Neural Collapse to propose a novel OOD detector. Specifically, we study the phenomenon that ID features tend to form clusters whereas OOD features reside far away. Inspired by Neural Collapse, we hypothesize and validate that ID features tend to cluster near weight vectors. We also explain why ID features tend to reside further from the origin and complement our method from this perspective. Experiments show that our method can achieve superior performance with low latency across diverse setups, improving upon the generalizability of existing work. We hope our work can inspire future work to explore the interplay between features and weight vectors for OOD detection and other research problems such as calibration and adversarial robustness.
Limitations:
Despite significantly reducing the generalization discrepancy, our $\mathtt{NCI}$ does not completely close the gap. Particularly, there is still an absolute difference of 1.72% in AUROC compared to the best result on the ImageNet ResNet-50 benchmark.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in our abstract and introduction are justified in rest of our paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss our limitations after the conclusion session.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We include complete proof and assumptions for our theoretical result in Section 3.1.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include code to reproduce our result in supplemental material.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We include our code in supplemental material and all data we used are open sourced.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify our train/test split and hyperparameter choice in our Experiment session.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We follow prior work and adhere to standard benchmarks with fixed train/test splits and pre-trained models, resulting in minimal randomness.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We use off-the-shelf models with references given in appendix. We also report the computer resources used for inference.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We reviewed and followed the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The work can enhance the safe deployment of AI models. We do not foresee any negative societal impact of this work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly credited the code, data, and models usded in the paper.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release our well-documented code alongside the paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
Appendix A Implementation Details
A.1 CIFAR-10
ResNet-18 For visualization in Fig. 1Left, Middle, we use a CIFAR-10 classifier of ResNet-18 backbone trained with cross-entropy loss. The classifier is trained for 100 epochs, with the initial learning rate 0.1 decaying to 0.01, 0.001, and 0.0001 at epochs 50, 75, and 90 respectively. For experiments in Table 1(a), we use the pre-trained model provided by the OpenOOD benchmark. And we refer readers to Zhang et al. (2023) for their training recipe.
DenseNet-101 For experiments on CIFAR-10 Benchmark presented in Table 6, we evaluate a CIFAR-10 classifier of DenseNet-101 backbone. The classifier is trained following the setups in Huang et al. (2017) with depth $L=100$ and growth rate $k=12$.
A.2 CIFAR-100
DenseNet-101 For experiments on the CIFAR-100 Benchmark presented in Table 6, we evaluate a CIFAR-100 classifier of the DenseNet-101 backbone. The classifier is trained following the setups in Huang et al. (2017) with depth $L=100$ and growth rate $k=12$.
A.3 ImageNet
@@
Swin v2 In Table 2, we use the $\mathtt{timm}$ (Wightman, 2019) implementation of Swin v2 as well as their pre-trained checkpoint ’swinv2_base_window8_256’.
Appendix B Alternatives Proximity Metrics
@@ Appendix C Baseline Methods
MSP Hendrycks and Gimpel (2016) proposes to detect OOD based on the maximum softmax probability. Given the penultimate feature $\bm{h}$ for a given test sample $\bm{x}$, the detection score of MSP can be represented as:
@@
ODIN Liang et al. (2018) proposes to amplify ID the OOD separation on top of MSP through temperature scaling and adversarial perturbation. Given a sample $\bm{x}$, ODIN constructs a noisy sample $\bm{x^{\prime}}$ from $\bm{x}$. Denote the penultimate feature of the noisy sample $\bm{x^{\prime}}$ as $\bm{h^{\prime}}$, ODIN assigns OOD score following:
@@
Energy Liu et al. (2020) designs an energy-based score function over the logit output. Given a test sample $\bm{x}$ as well as its penultimate layer feature $\bm{h}$, the energy based detection score can be represented as:
@@
ReAct Sun et al. (2021) builds upon the energy score proposed in Liu et al. (2020) and regularizes the score by truncating the penultimate layer estimation. We set the truncation threshold at $90$ percentile in our experiments.
Dice Sun and Li (2022) builds upon the energy score proposed in Liu et al. (2020). Leveraging the observation that units and weights are used sparsely in ID inference, Sun and Li (2022) proposes to select and compute the energy score over a selected subset of weights based on their importance. We set a threshold at $90$ percentile for CIFAR experiments and $70$ percentile for ImageNet experiments following Sun and Li (2022).
ASH Djurisic et al. (2022) builds upon the energy score proposed in Liu et al. (2020). Prior to the Energy score, ASH sorts each feature to find the top-k elements, scales up the top-k elements, and sets the rest to zero. We note that in addition to the cost of Energy, ASH introduces a sorting cost of $O(P\log k)$, where $P$ is the penultimate layer dimension.
Scale Xu et al. (2023) builds upon the energy score proposed in Liu et al. (2020). Prior to the Energy score, Scale sorts each feature to find the top-k elements and based on the statistics, scales all elements in the feature. We note that in addition to the cost of Energy, Scale also introduces a sorting cost of $O(P\log k)$, where $P$ is the penultimate layer dimension.
Mahalanobis On the feature space, Lee et al. (2018) models the ID feature distribution as multivariate Gaussian and designs a Mahalanobis distance-based score:
@@
On top of the basic score, Lee et al. (2018) also proposes two techniques to enhance the OOD detection performance. The first is to inject noise into samples. The second is to learn a logistic regressor to combine scores across layers. We tune the noise magnitude and learn the logistic regressor on an adversarial constructed OOD dataset. The selected noise magnitude is 0.005 in both our ResNet and DenseNet experiments.
KNN Chen et al. (2020) proposes to detect OOD based on the k-th nearest neighbor distance between the normalized embedding of the test sample $\bm{z_{x}}/|\bm{z_{x}}|$ and the normalized training embeddings on the penultimate space. Chen et al. (2020) also observes that contrastive learning helps in improving OOD detection effectiveness.
GradNorm Huang et al. (2021) extracts information from the gradient space to detect OOD samples. Specifically, Huang et al. (2021) defines the OOD score function as the $\mathtt{L1}$ norm of the gradient of the weight matrix with respect to the KL divergence between the softmax prediction for $\bm{x}$ and the uniform distribution.
@@
ViM Wang et al. (2022) proposes to integrate class-specific information into feature space information by adding energy score to the feature norm in the residual space of the training feature matrix. The detection score is designed to be:
@@
fDBD Liu and Qin (2023) proposes to detect OOD based on estimated feature distance to decision boundaries of class $c\in\mathcal{C}$ besides its predicted class $f(\bm{x})$:
@@
Appendix D Evaluation on DenseNet

### v7

@@ 5 Related Work
OOD Detection
Extensive research has focused on OOD detection algorithms. One line of work is post-hoc and builds upon pre-trained models. For example, [13, 21, 24, 38, 37, 22, 46] design OOD score over the output space of a classifier. Meanwhile, [20] and [39] measure OOD-ness from the perspective of ID clustering in feature space. Our work extends the observation that ID features tend to cluster from the perspective of Neural Collapse. While existing work is more focused on certain classification tasks, our proposed OOD detector is tested to be highly versatile.
Others [36, 34, 35, 52] explore the regularization of OOD detection in training. For example, [7, 14] propose OOD-specific architecture whereas [16, 44] design OOD-specific training loss. [40] brings attention to representation learning for OOD detection and proposes an OOD-specific contrastive learning scheme. Our work does not belong to this school of thought and is not restricted to specific training schemes or architecture. Notably, as shown in Appendix 10, our NCI can also benefit from training-time algorithms.
Neural Collapse
Neural Collapse was first observed in [32]. During Neural Collapse, the penultimate layer features collapse to class means, the class means and the classifier collapses to a simplex equiangular tight framework, and the classifier simplifies to adopt the nearest class-mean decision rule. Further work has provided theoretical justification for the emergence of Neural Collapse [10, 30, 51, 53]. Application-wise, [53] derives an efficient training algorithm inspired by Neural Collapse. Our concurrent work [1] also leverages Neural Collapse for OOD detection but overlooks class-specific information revealed by Neural Collapse, which is central to our approach.
@@ 6 Conclusion
This work leverages insights from Neural Collapse to propose a novel OOD detector. Specifically, we study the phenomenon that ID features tend to form clusters whereas OOD features reside far away. Inspired by the trend of Neural Collapse prevalent on practical models, we hypothesize and validate that ID features tend to cluster near weight vectors. We also explain why ID features tend to reside further from the origin and complement our method from this perspective. Experiments show the effectiveness of our method on practical models without requiring the complete convergence of Neural Collapse. Further, our method improves the overall performance with minimal latency across diverse benchmarks. We hope our work can inspire future work to explore the interplay between features and weight vectors for OOD detection and other research problems such as calibration and adversarial robustness.
Acknowledgments
We would like to express our sincere gratitude to Yubing Jian for conducting several experiments that made significant contributions to this work and for his valuable discussions.
Supplementary Material
7 Implementation Details
7.1 CIFAR-10
ResNet-18 For visualization in Fig. 2Left, Middle, we use a CIFAR-10 classifier of ResNet-18 backbone trained with cross-entropy loss. The classifier is trained for 100 epochs, with the initial learning rate 0.1 decaying to 0.01, 0.001, and 0.0001 at epochs 50, 75, and 90 respectively. For experiments in Table 1(a), we use the pre-trained model provided by the OpenOOD benchmark. And we refer readers to [49] for their training recipe.
DenseNet-101 For experiments on CIFAR-10 Benchmark presented in Table 9, we evaluate a CIFAR-10 classifier of DenseNet-101 backbone. The classifier is trained following the setups in [15] with depth $L=100$ and growth rate $k=12$.
7.2 CIFAR-100
DenseNet-101 For experiments on the CIFAR-100 Benchmark presented in Table 9, we evaluate a CIFAR-100 classifier of the DenseNet-101 backbone. The classifier is trained following the setups in [15] with depth $L=100$ and growth rate $k=12$.
7.3 ImageNet
@@
Swin v2 In Table 2, we use the $\mathtt{timm}$ [45] implementation of Swin v2 as well as their pre-trained checkpoint ’swinv2_base_window8_256’.
8 Alternatives Proximity Metrics
@@ 9 Baseline Methods
MSP [12] proposes to detect OOD based on the maximum softmax probability. Given the penultimate feature $\bm{h}$ for a given test sample $\bm{x}$, the detection score of MSP can be represented as:
@@
ODIN [21] proposes to amplify ID the OOD separation on top of MSP through temperature scaling and adversarial perturbation. Given a sample $\bm{x}$, ODIN constructs a noisy sample $\bm{x^{\prime}}$ from $\bm{x}$. Denote the penultimate feature of the noisy sample $\bm{x^{\prime}}$ as $\bm{h^{\prime}}$, ODIN assigns OOD score following:
@@
Energy [24] designs an energy-based score function over the logit output. Given a test sample $\bm{x}$ as well as its penultimate layer feature $\bm{h}$, the energy based detection score can be represented as:
@@
ReAct [38] builds upon the energy score proposed in [24] and regularizes the score by truncating the penultimate layer estimation. We set the truncation threshold at $90$ percentile in our experiments.
Dice [37] builds upon the energy score proposed in [24]. Leveraging the observation that units and weights are used sparsely in ID inference, [37] proposes to select and compute the energy score over a selected subset of weights based on their importance. We set a threshold at $90$ percentile for CIFAR experiments and $70$ percentile for ImageNet experiments following [37].
ASH [8] builds upon the energy score proposed in [24]. Prior to the Energy score, ASH sorts each feature to find the top-k elements, scales up the top-k elements, and sets the rest to zero. We note that in addition to the cost of Energy, ASH introduces a sorting cost of $O(P\log k)$, where $P$ is the penultimate layer dimension.
Scale [47] builds upon the energy score proposed in [24]. Prior to the Energy score, Scale sorts each feature to find the top-k elements and based on the statistics, scales all elements in the feature. We note that in addition to the cost of Energy, Scale also introduces a sorting cost of $O(P\log k)$, where $P$ is the penultimate layer dimension.
Mahalanobis On the feature space, [20] models the ID feature distribution as multivariate Gaussian and designs a Mahalanobis distance-based score:
@@
On top of the basic score, [20] also proposes two techniques to enhance the OOD detection performance. The first is to inject noise into samples. The second is to learn a logistic regressor to combine scores across layers. We tune the noise magnitude and learn the logistic regressor on an adversarial constructed OOD dataset. The selected noise magnitude is 0.005 in both our ResNet and DenseNet experiments.
KNN [3] proposes to detect OOD based on the k-th nearest neighbor distance between the normalized embedding of the test sample $\bm{z_{x}}/|\bm{z_{x}}|$ and the normalized training embeddings on the penultimate space. [3] also observes that contrastive learning helps in improving OOD detection effectiveness.
GradNorm [17] extracts information from the gradient space to detect OOD samples. Specifically, [17] defines the OOD score function as the $\mathtt{L1}$ norm of the gradient of the weight matrix with respect to the KL divergence between the softmax prediction for $\bm{x}$ and the uniform distribution.
@@
ViM [43] proposes to integrate class-specific information into feature space information by adding energy score to the feature norm in the residual space of the training feature matrix. The detection score is designed to be:
@@
fDBD [22] proposes to detect OOD based on estimated feature distance to decision boundaries of class $c\in\mathcal{C}$ besides its predicted class $f(\bm{x})$:
@@
10 Performance Boosting with Training-Time Regularization
In this section, we investigate the compatibility of NCI with training-time regularization algorithms for OOD detection. On the CIFAR-10 benchmark, we evaluate NCI performance on models trained with a training-time regularization method, T2FNorm[35], as an example. We also compare this with NCI performance on models trained using standard cross-entropy loss, as presented in Table 1. In Table 8, we report the average AUROC across near OOD datasets (CIFAR-100 and TIN) and far OOD datasets (MNIST, SVHN, Texture, and Place365). We observe a performance boost in both cases, demonstrating the compatibility of NCI as a post-hoc method with training-time regularization algorithms. This also highlights the effectiveness of combining NCI with training-time regularization for improved OOD detection performance.
11 Additional Baselines
In the setup of Section 4.1, we further compare our NCI with two additional baselines, GEN [25] and SHE [48]. On the CIFAR-10 benchmark, SHE, GEN, and NCI achieve average AUROCs of 84.06, 90.30, and 90.46, respectively. On the ImageNet benchmark, SHE, GEN, and NCI achieve average AUROCs of 84.06, 86.20, and 88.56. Our result further demonstrates NCI’s superior performance in mitigating generalization dependency across different classification tasks.
In the setup of Section 4.2, we apply the CLIP-based method MCM [29] (for vision-language models) to the vision models studied in that section. For MCM, we report the best average AUROC scores achieved through temperature sweeping across [0.01,0.1,1,10] for each model. On ViT, MCM achieves an average AUROC of 93.60, whereas NCI achieves 93.64. On Swin v2, MCM achieves an average AUROC of 77.73, whereas NCI achieves 80.94, demonstrating NCI’s superior performance in this setup.
Table 8: Performance Boosting with Training-Time Regularization. On the CIFAR-10 benchmark, we report the average AUROC for Near and Far OOD. Experiments are conducted with ResNet-18 trained using Cross-Entropy (CE) and T2FNorm.
12 Evaluation on DenseNet